Please help me with this: Find (f-g)(5) when f(x)=5x square - 7 and g(x)= x + 2
Thank you

Answer by stanbon(75887)

Find (f-g)(5) when f(x)=5x square - 7 and g(x)= x + 2
---
(f-g)(5) = f(5)-g(5)
---
= [5(5^2)-7]- [5+2]
---
= [125-7]-7
---
= 125-14
= 111
===========
Cheers,
Stan H.
===========